OpenStudy (anonymous):

Factor 30b^2 + 48b – 24 (1 point) 24(6b^2 + 2b – 1) 6(5b – 2)(b + 2) 4(5b – 2)(b + 2) 6(5b + 2)(b + 2)

hartnn (hartnn):

u already know the answer, isn't it ? just want to confirm?

OpenStudy (anonymous):

actually no not really..i want someone to walk me throughthe steps but not tell me the answer...i actually dont get this one AT ALL

hartnn (hartnn):

what can u see common between 30,48 and 24 ?

OpenStudy (anonymous):

6

hartnn (hartnn):

so take that common out, what remains inside brackets?

OpenStudy (anonymous):

i do not understand what you are saying but I no you can already delete a and c..i dont get when you say take sommon out

hartnn (hartnn):

6 multiplied by what is 30?

OpenStudy (anonymous):

5

hartnn (hartnn):

what about 48 and 24?

OpenStudy (anonymous):

6 times 8 equals 48 and 6 times 4 equals 24

hartnn (hartnn):

so u get 6(5b^2+8b-4) did u get this step?

OpenStudy (anonymous):

yea

hartnn (hartnn):

now, tell me 2 numbers whose product is 5*-4 = -20 and whose sum is 8

OpenStudy (anonymous):

5^5-4=-20

hartnn (hartnn):

what!! the numbers are 10 and -2.....

OpenStudy (anonymous):

how 5^10-4 do not equal -20

hartnn (hartnn):

i never asked u to do ^ operation... product =-20 .... 10*-2 = -20 sum = 8.......10-2 = 8

OpenStudy (anonymous):

Calm down..im sorry i did not get it .. imma just move onto the next question..just forget it